OpenStudy (baby456):

3. Line A passes through points (2,4) and (3,6).Write an equation of a line that is parallel to line A.Wrie an equation of a line that will be perpendicular to line A.

OpenStudy (baby456):

@jdoe0001

OpenStudy (baby456):

@jim_thompson5910

OpenStudy (baby456):

@zepdrix

OpenStudy (jdoe0001):

a line parallel to line A will have the same slope as line A thus, find the slope of line A then use that, and slap it on an equation in slope-intercept form with whatever y-intercept you wish

OpenStudy (baby456):

the slope of line a is 2/1 @jdoe0001

OpenStudy (jdoe0001):

yeap is 2/1 or 2 so just slap it on any slope-intercept form say y - mx + b m = 2 b = whatever, say hmm 17 so \(\large \bf y = 2x+17\) it has the same slope as A, thus is parallel

OpenStudy (jdoe0001):

now the perpendicular well, we know A slope is 2 a line perpendicular to it will have a slope of, NEGATIVE RECIPROCAL meaning \(\bf slope=\cfrac{2}{{\color{blue}{ 1}}}\qquad negative\implies -\cfrac{2}{{\color{blue}{ 1}}}\qquad reciprocal\implies - \cfrac{{\color{blue}{ 1}}}{2}\) so now we know the slope of a line perpendicular to A is -1/2 same slap it on a slope-intercept form m = -1/2 b = whatever... say hmmm 55 thus \(\large \bf y=-\cfrac{1}{2}x+55\)